# Euclid preparation

### v1

@@
Euclid is a new space mission within the European Space Agency’s (ESA) Cosmic Vision 2015–2025 programme, aiming to study the nature of dark matter and dark energy by estimating the expansion history of the Universe and the growth rate of cosmic structures (Laureijs et al., 2011; Euclid Collaboration: Mellier et al., 2025). For this purpose Euclid was designed as a wide-field telescope in the visible to near-infrared (NIR) wavelength range to survey 14 000 deg2 of extragalactic sky from the Sun–Earth Lagrange point 2 (Euclid Collaboration: Scaramella et al., 2022).
@@
To achieve its objectives, Euclid’s payload consists of two scientific instruments developed by the Euclid Consortium: VIS provides high-resolution images of the sky in a single broad wavelength band from 550 nm to 950 nm (Euclid Collaboration: Cropper et al., 2025). With a spatial sampling of and a tight control on the point spread function (PSF), the VIS instrument is optimized for the the measurement of galaxy shapes for weak lensing, which enables the mapping of dark matter structure as tool for cosmological studies (Cropper et al., 2016; Euclid Collaboration: Cropper et al., 2025). Euclid’s second instrument is the Near-Infrared Spectrometer and Photometer (NISP) operating in the NIR (Euclid Collaboration: Jahnke et al., 2025), sharing with VIS a common field-of-view (FoV) of 0.54 deg2 (Euclid Collaboration: Mellier et al., 2025). NISP has been designed for the spectroscopic measurements of galaxy distances for the galaxy clustering probe, as well as NIR information for photometric redshifts.
@@
As a cosmology mission, Euclid’s ultimate goal is to reach a combined cosmological ‘Figure of Merit’ FoM 400 (Euclid Collaboration: Mellier et al., 2025) on the dark energy parameters with weak lensing and galaxy clustering. Reaching this ambitious goal requires a combined approach that leverages both a comprehensive joint analysis of weak gravitational lensing, galaxy-galaxy lensing, and galaxy clustering (commonly referred to as a -point analysis) and a three-dimensional galaxy clustering analysis using spectroscopic redshifts to evaluate cosmological distances. To that end, the Euclid spectroscopic survey is designed to target H line-emitteing galaxies in the redshift range 0.9–1.8, thereby providing precise measurements of the growth of structure and redshift-space distortions as well as the expansion history via baryon acoustic oscillations. Consequently, the galaxy clustering analyses requires to obtain spectroscopic redshift for at least 1700 galaxies per square degree. This corresponds to redshift measurements for 2.55107 galaxies over the survey area of 14 000 deg2 of the Euclid Wide Survey (Euclid Collaboration: Scaramella et al., 2022) achieved through slit-less spectroscopy. The Euclid core science also requires the galaxy redshifts to be estimated with an accuracy , which translates to a minimal spectral resolution when considering a source size of .
@@
In the following we will describe the NISP spectroscopic channel that was designed and built for Euclid’s NIR spectroscopic survey. Section 2 provides a general overview of the NISP instrument as well as the grisms as the key components for NISP’s spectroscopic capabilities. The setup for NISP ground tests of the spectroscopic channel are described in Sect. 3. NISP’s spectroscopic optical quality is presented in Sect. 4 and Sect. 6 shows a detailed analysis of the grisms’ dispersion properties. In Sect. 7 the overall resulting spectroscopic performance is discussed. The paper closes with a brief conclusion and a description of available data products characterising the grisms in Sect. 8.
@@ 2.1 Overview of NISP components and properties
NISP consists of a Silicon Carbide (SiC) mechanical structure (NISP Structure Assembly STructure – NI-SA-ST) that houses the detector system (NI-DS) on one side and the optical system on the other (Bougoin et al., 2017). The SiC was chosen for its stiffness and its thermal stability allowing to maintain optical alignement and stability over the mission lifetime. Additionally, fine tuning of internal NISP heaters maintain the opto-mechanics temperature in the required range around – to further maintain optical alignment. Low conductance ( in total) bipods and monopod (NISP Structure Assembly HexaPods – NI-SA-HP) interface the NISP instrument with the the Euclid Payload Module (PLM) baseplate.
The NI-DS (Bonnefoi et al., 2016) acquires images by sampling the FoV with 16 HgCdTe 2k 2k pixels NIR detectors produced by Teledyne Imaging Sensors, arranged in a 4 4 mosaic. Their pixel size of 18 provides NISP with a spatial sampling of the sky of , and the detectors have their wavelength cut-off tuned to 2.3 to minimize thermal noise from the optical system’s thermal emission. Maintaining the focal plane below is crucial for optimal performance and is achieved through thermal coupling of the instrument with an external radiator using four highly conductive thermal straps. Furthermore, NISP’s detectors are read out by 16 individual SIDECAR ASICs, each operating at approximately and dissipating up to of heat. To manage this heat load effectively, high-efficiency thermal coupling to the baseplate is employed, facilitated by methane heat pipes (MHP)s. Due to the substantial heat dissipation, a large radiator of approximately is required to radiate this heat into space, accounting for both the electronics’ heat and any conductive or radiative leaks between the PLM and the radiator.
The NISP instrument images the sky in two different channels: a photometric channel for the acquisition of images with broadband filters and a spectrometric channel for the acquisition of slit-less dispersed images. The optical system of NISP (Bodendorf et al., 2019; Grupp et al., 2012) comprises several elements; a spherical-aspherical meniscus-type corrector lens (CoLA) at the entrance of NISP that actively takes part in correcting the chromatic aberration of the image; the filter wheel (NI-FWA) with a set of three filters (Euclid Collaboration: Schirmer et al., 2022), namely (950–1212 nm), (1168–1567 nm) and (1522–2021 nm) as well as a ‘closed’ position, blocking light from the telescope; the grism wheel (NI-GWA) with a set of three red grisms ( nm), a single blue grism ( nm) and an open position; and finally an assembly of three lenses which, together with their holding structure, constitute the camera lens assembly (CaLA), focussing light onto the detector plane.
@@ 2.2 The NISP grisms
The key components in creating dispersed light for determining redshifts of distant galaxies are the NISP grisms. They are the combination of a transmission grating with a prism. This combination allows us to create a common FoV with the NISP photometry channel, placing the dispersed light onto the same detector plane. The NISP instrument contains four different grisms labelled RGS000, RGS180, RGS270, BGS000, three in a red and one in a blue bandpass.
Each of the grisms comprises several optical elements, each fulfilling a specific function (Costille et al., 2019), presented in Fig. 2. A resin-free blazed dispersion grating (dark blue) is directly engraved onto the hypotenuse surface of a fuse-silica prism (light blue) to disperse the light. Each grism has a diameter of 140 mm and a central thickness of approximately 12 mm (Costille et al., 2014), making them the largest grisms ever flown into space. The prism and the grating design were defined to maximize transmission efficiency at the un-deflected wavelength of about 1.2 for the red and 0.9 for the blue grisms, respectively. Additionally, the grating groove has been engraved with a curvature to apply a spectral wavefront correction to the incoming light. The base of the prism carries optical power, that is, it has a curved surface (dashed line) to provide a fine-focus correction for the incoming light beam towards the NISP detection plane. Finally, a multilayer coating (yellow) has been deposited on the curved surface to define each grism’s transmission bandpass.
@@
The grisms create different orientations of spectral dispersion on the focal plane. With a view from the NISP optics towards the projected spectra on the detector plane and with grism RGS000 as the reference, the grism RGS180 has a 180∘ orientation and hence disperses light in the opposite direction as RGS000. Similarly, RGS270 has a 270∘ counter-clockwise orientation with respect to RGS000. These three different orientation were chosen to permit ‘decontamination’ of spectra from overlaps with other sources, creating a reconstructed spectrum from successively acquired exposures of a given field with the three red grisms.
Each grism is glued into a mechanical mount that provides the mechanical interface of the grism with the wheel structure, in order to both provide a defined position as well as minimise stresses from cool-down to the cryogenic environment of the instrument during operations.
@@ 3 NISP ground testing setup
After completion of the instrument integration in 2019, the NISP flight model underwent a series of tests in vacuum and with the instrument cooled down to its operational temperature. Two test campaigns were conducted inside the ERIOS vacuum chamber at the ‘Laboratoire d’Astrophysique de Marseille’ (LAM) in the fall of 2019 and the beginning of 2020, to validate the instrument’s performance and its functionality in a space environment.
The ERIOS chamber (Costille et al., 2016b) is a large cryogenic chamber, with an external envelope of 6 m in length and a diameter of 4 m, capable to achieve high vacuum ( 10-6 mbar) and to cool down an inner volume of 45 m3 to low temperatures ( 80 K) with a high stability (4 mK) thanks to liquid nitrogen shrouds covering the entire inner surface of the chamber.
@@
The entire aperture of the mirror was illuminated through this 2 pin-hole located at the primary focus of the elliptical mirror, creating a point source for NISP. Both, the mirror and the pin-hole were attached to the same baseplate and the whole system (mirror, pin-hole, and baseplate) was made from silica to ensure stability of the mirror’s focal distance. The telescope baseplate was attached to translation and rotation stages using thermal flexures made from a low conductivity material, thermally isolating the telescope, while providing it with 5 degrees of freedom when being pointed towards different positions in NISP’s FoV.
@@
Inside ERIOS, light was fed to the telescope through a cryogenic fibre connected to the pin-hole, with light sources being located outside the chamber. At the interface of the ERIOS chamber a pair of vacuum windows connected the fibre with the light sources, using a cooled neutral density filter to suppress thermal background. On the outside, a set of neutral density filters allowed to select the intensity of light entering the fibre.
@@
Spectral lamps were fed into the monochromator to characterise and calibrate it, using HgCd, He, or Cs lamps for the visible band below , and an Ar lamp for the near-infrared domain. The output of the monochromator was scanned using either silicon or germanium photodiodes. The calibration yielded a selectable maximum wavelength accuracy of , and the monochromator bandwidth was measured to be approximately in Full-Width-Half-Maximum (FWHM).
The etalon was re-aligned and re-calibrated in warm conditions with a Perkin Elmer Lambda 900 UV/VIS/NIR spectrometer (‘Lambda 900’) before testing of NISP started. During the tests, thermal sensor monitored the stability of the light sources and the etalon. Figure 3 shows the Fabry–Pérot etalon’s spectrum as it was measured and calibrated with a precision of 0.2 nm using the Lambda 900 spectrometer. A comparison to the grisms’s transmission shows the number of Fabry–Pérot’s peaks available for calibration. The grism channels’ total transmission shown here accounts for the transmission of the CoLA and CaLA lenses, the grisms themselves, as well as the detector quantum efficiency. In this figure neither the Euclid mirrors nor dichroic are accounted for, as both elements were not present in the instrument-level test setup. Note that in Fig. 3, the transmission of the -order components of NISP’s spectrograms, which are below few percent at maximum, have been multiplied by a factor of 10 to make them stand out.
Prior to the start of the optical tests, the focal length of the NISP instrument was measured at cold (NISP focal plane at and NISP optics at ), using a monochromatic light beam with a wavelength of 1000 nm in the photometric passband. In this process, PSF widths were measured from spots created by the point source. This was carried out at five positions in the NISP FoV, where hundreds of individual slightly dithered exposures were taken to over-sample the PSF and to increase statistical accuracy on the PSF width measurement. This was repeated for different distances of the telescope simulator from NISP. During the PSFs acquisitions, the relative distance between NISP and the telescope simulator was accurately monitored with to a laser tracker system that measured the position of ten targets attached to the NISP mechanical structure relative to five targets attached to the back of the telescope mirror (Costille et al., 2016a, 2018).
@@ 4.1 Optical quality of RGS000, RGS180 & BGS000
The NISP optical quality was assessed by evaluating the radii that encircle 50% and 80% of the PSF’s total energy – in the following these radii will be referred to as EE50 and EE80, respectively. In the spectroscopic channel, this verification was done for every grism of the NI-FWA at three monochromatic wavelengths: 1300 nm, 1500 nm, and 1800 nm for all three red grisms and 900 nm, 1120 nm, and 1340 nm for the blue grism. Additionally, we tested image quality of both RGS000 and RGS180 with a grism wheel rotation offset to validate that offsetting the grism positions by this amount would not degrade image quality. This additional measurement was part of an evaluation of a modified survey strategy to overcome the reduced quality of data from RGS270 (see Sect. 5).
@@
The image quality was evaluated with an analysis of monochromatic PSF images taken at the four corners of the NISP FoV. Because the NISP plate scale of per pixel undersamples the PSF, hundreds of PSF images were acquired around each position, dithering by a tenth of a pixel around the target position in the NISP image plane. This dithering, obtained by moving the telescope on its axes, allowed to spatially over-sample the PSFs and reduced the statistical uncertainties on the EE50 and EE80 measurements.
@@
Before measuring PSF properties, images acquired by the NISP instrument were corrected for pixel quantum efficiency and conversion gain, and a NISP residual thermal background acquired under ‘dark test conditions’, i.e. with the NISP NI-FWA in closed position and all light sources turned off. An automatic procedure, derived from the work of Jiang et al. (2016), was used to locate the PSF within the 2k 2k pixels of the target detectors. Once located, a 20 20 pixel stamp centred on the expected PSF position was extracted from the images and PSF metrics were extracted by image analyses. Every extracted sub-image was controlled by eye and e.g. detector cosmetic defects misidentified as PSFs by our algorithm were rejected from the analyses. An Intra Pixel Capacitance (IPC) correction, relying on a Richardson–Lucy deconvolution method (Lucy, 1974; Richardson, 1972) using IPC coefficients previously measured during detector characterization as the deconvolution kernel, was applied to each extracted stamp. However, it should be noted that the spread of the measured EE50 and EE80 distributions were much wider than the correction factor applied through this IPC deconvolution. The IPC signal leaking to neighbouring pixels was measured to be very small, of 2.870.01% losses in the central pixel that are non-equally redistributed to the eight immediately adjacent pixels (Le Graët et al., 2022).
@@
Figure 4 shows one example of a NISP PSF acquisition. The left panel is a full frame image (i.e. 2k 2k pixels) from one detector, and the artificial point source at the top. A ring-like structure is visible surrounding the PSF, that resulted from scattered light at the exit of the pin-hole. Since the ring radius is very large compared to the PSF diameter, it was easily disregarded in the analyses. However, these rings were quite useful in validating that the blind search algorithm indeed successfully located the PSF. In the centre panel of Fig. 4 a zoom into the small box around the PSF is shown. The diffuse circular feature to the right of the compact PSF is produced by light emission coming from the fibre core passing through a non-perfectly blocking neutral density filter around the pin-hole. This was meant to block light throughout the full NISP wavelength range. It however appeared that instead it had a transmission increasing with wavelength, starting with an optical density 5 at 900 nm and reaching an optical density of 4.5 at 1800 nm, creating this patch. The rightmost panel shows another zoom step into the PSF, with contrast adjusted to prevent display saturation.
@@
The EE50 and EE80 are deduced from PSF model fitting with the following model
@@
where are the spatial coordinates in the pixel mosaic reference frame, is a PSF model of parameters – for the functions that were considered see below – is a -Bell function (Boyd, 2006) of parameters – which is used to model the contribution from the fibre core – and , a constant introduced to account for a residual constant background. As the fit is limited to a small stamp of 20 20 pixels centred on the PSF, a constant background is a good first approximation at these scales. Note that the axis is defined such that the light dispersed by RGS000 along a spectral trace has its wavelength increase with .
To estimate the reliability of our estimates, we tested three different functions to model the PSF profile.
@@
An asymmetric-Gaussian function for which amplitude, centroid position, width along the and axes as well as inclination with respect to the axis are free parameters. When using this model, EE50 is evaluated as , with the width of the PSF in the cross-dispersion direction. We used the cross-dispersion direction to minimise the impact of the monochromator bandwidth on the estimate of EE50. However, such an estimation might be underestimating the true width, by assuming the PSF is having a Gaussian profile and neglecting potential contributions from a non-Gaussian tail.
@@
The sum of two asymmetric-Gaussian functions, hereafter refereed as a dual-Gaussian profile, for which the respective amplitude, common centroid position, individual width along the and axes, as well as common inclination are free parameters. To avoid any degeneracy in the model fitting, the amplitude and width of the larger Gaussian was defined by a multiplicative factor relative to that of the narrower Gaussian. This model tends to better account for potential non-Gaussian tails in the PSF profile.

### v2

@@
Euclid is a new space mission within the European Space Agency’s (ESA) Cosmic Vision 2015–2025 programme, aiming to study the nature of dark matter and dark energy by estimating the expansion history of the Universe and the growth rate of cosmic structures (Laureijs et al., 2011; Euclid Collaboration: Mellier et al., 2025). For this purpose Euclid was designed as a wide-field telescope in the visible to near-infrared wavelength range to survey 14 000 deg2 of extragalactic sky from the Sun–Earth Lagrange point 2 (Euclid Collaboration: Scaramella et al., 2022).
@@
To achieve its objectives, Euclid’s payload consists of two scientific instruments developed by the Euclid Consortium: VIS provides high-resolution images of the sky in a single broad wavelength band from 550 nm to 950 nm (Euclid Collaboration: Cropper et al., 2025). With a spatial sampling of and a tight control on the point spread function, the VIS instrument is optimised for the measurement of galaxy shapes for weak lensing, which enables the mapping of dark matter structure as a tool for cosmological studies (Cropper et al., 2016; Euclid Collaboration: Cropper et al., 2025). Euclid’s second instrument, the Near-Infrared Spectrometer and Photometer (Euclid Collaboration: Jahnke et al., 2025), operates in the near-infrared and shares a common field-of-view of 0.54 deg2 with VIS (Euclid Collaboration: Mellier et al., 2025) thanks to a dichroic transparent to near-infrared which is located in front of the NISP entrance pupil. Near-Infrared Spectrometer and Photometer has been designed for the spectroscopic measurements of galaxy distances for the galaxy clustering probe, as well as near-infrared information for photometric redshifts.
@@
As a cosmology mission, Euclid’s ultimate goal is to reach a combined cosmological ‘Figure of Merit’ FoM 400 (Euclid Collaboration: Mellier et al., 2025) on the dark energy parameters with weak lensing and galaxy clustering. Reaching this ambitious goal requires a combined approach that leverages both a comprehensive joint analysis of weak gravitational lensing, galaxy-galaxy lensing, and galaxy clustering (commonly referred to as a -point analysis) and a three-dimensional galaxy clustering analysis using spectroscopic redshifts to evaluate cosmological distances. To that end, the Euclid spectroscopic survey is designed to target H line-emitting galaxies in the redshift range 0.9–1.8, thereby providing precise measurements of the growth of structure and redshift-space distortions as well as the expansion history via baryon acoustic oscillations. Consequently, the galaxy clustering analyses requires to obtain spectroscopic redshift for at least 1700 galaxies per square degree. This corresponds to redshift measurements for 2.55107 galaxies over the survey area of 14 000 deg2 of the Euclid Wide Survey (Euclid Collaboration: Scaramella et al., 2022) achieved through slitless spectroscopy. The Euclid core science also requires the galaxy redshifts to be estimated with an accuracy , which translates to a minimal spectral resolution when considering a source size of .
@@
This work focuses on evaluating NISP spectroscopic performance through the analysis of point spread function quality and spectral dispersion properties. In the following we will describe the NISP spectroscopic channel that was designed and built for Euclid’s near-infrared spectroscopic survey. Section 2 provides a general overview of the Near-Infrared Spectrometer and Photometer instrument as well as the grisms as the key components for NISP’s spectroscopic capabilities. The setup for NISP ground tests of the spectroscopic channel are described in Sect. 3. NISP’s spectroscopic optical quality is presented in Sect. 4 and Sect. 5 shows a detailed analysis of the grisms’ dispersion properties. In Sect. 6 the overall resulting spectroscopic performance is discussed. The paper closes with a brief conclusion and a description of available data products characterising the grisms in Sect. 7.
@@ 2.1 Overview of NISP components and properties
NISP consists of a Silicon Carbide mechanical structure (NISP Structure Assembly STructure – NI-SA-ST) that houses the detector system on one side and the optical system on the other (Bougoin et al., 2017). The Silicon Carbide was chosen for its stiffness and its thermal stability allowing to maintain optical alignment and stability over the mission lifetime. Additionally, fine tuning of internal NISP heaters maintain the opto-mechanics temperature in the required range around – to further maintain optical alignment. Low conductance ( in total) bipods and monopod (NISP Structure Assembly HexaPods – NI-SA-HP) interface the NISP instrument with the Euclid Payload Module baseplate.
The detector system (Bonnefoi et al., 2016) acquires images by sampling the field-of-view with 16 HgCdTe 2k 2k pixels near-infrared detectors produced by Teledyne Imaging Sensors, arranged in a 4 4 mosaic. Their pixel size of 18 provides NISP with a spatial sampling of the sky of , and the detectors have their wavelength cut-off tuned to 2.3 to minimise thermal noise from the optical system’s thermal emission. Maintaining the focal plane below is crucial for optimal performance and is achieved through thermal coupling of the instrument with an external radiator using four highly conductive thermal straps. Furthermore, NISP’s detectors are read out by 16 individual SIDECAR ASICs, each operating at approximately and dissipating up to of heat. To manage this heat load effectively, high-efficiency thermal coupling to the baseplate is employed, facilitated by methane heat pipess. Due to the substantial heat dissipation, a large radiator of approximately is required to radiate this heat into space, accounting for both the electronics’ heat and any conductive or radiative leaks between the Payload Module and the radiator.
The NISP instrument images the sky in two different channels: a photometric channel for the acquisition of images with broadband filters and a spectrometric channel for the acquisition of slitless dispersed images. The optical system of NISP (Bodendorf et al., 2019; Grupp et al., 2012) comprises several elements; a spherical-aspherical meniscus-type corrector lens at the entrance of NISP that actively takes part in correcting the chromatic aberration of the image; the filter wheel with a set of three filters (Euclid Collaboration: Schirmer et al., 2022), namely (950–1212 nm), (1168–1567 nm) and (1522–2021 nm) as well as a ‘closed’ position, blocking light from the telescope; the grism wheel with a set of three red grisms ( nm), a single blue grism ( nm) and an open position; and finally an assembly of three lenses which, together with their holding structure, constitute the camera lens assembly, focusing light onto the detector plane.
@@ 2.2 The NISP grisms
The key components in creating dispersed light for determining redshifts of distant galaxies are the NISP grisms. They are the combination of a transmission grating with a prism. This combination allows us to create a common field-of-view with the NISP photometry channel, placing the dispersed light onto the same detector plane. The NISP instrument contains four different grisms labelled RGS000, RGS180, RGS270, BGS000, three in a red and one in a blue bandpass.
Each of the grisms comprises several optical elements, each fulfilling a specific function (Costille et al., 2019), presented in Fig. 2. A resin-free blazed dispersion grating (dark blue) is directly engraved onto the hypotenuse surface of a fused-silica prism (light blue) to disperse the light. Each grism has a diameter of 140 mm and a central thickness of approximately 12 mm (Costille et al., 2014), making them the largest grisms ever flown into space. The prism and the grating design were defined to maximise transmission efficiency at the undeflected wavelength of about 1.2 for the red and 0.9 for the blue grisms, with transmission required to be higher than in grism bandpass. Additionally, the grating groove has been engraved with a curvature on the hypotenuse surface of the prism to apply a spectral wavefront correction to the incoming light. The base of the prism carries optical power, that is, it has a curved surface (dashed line) to provide a fine-focus correction for the incoming light beam towards the NISP detection plane. Finally, a multilayer coating (yellow) has been deposited on the curved surface to define each grism’s transmission bandpass.
@@
The grisms create different orientations of spectral dispersion on the focal plane. With a view from the NISP optics towards the projected spectra on the detector plane and with grism RGS000 as the reference, the grism RGS180 has a 180∘ orientation and hence disperses light in the opposite direction to RGS000. Similarly, RGS270 has a 270∘ counter-clockwise orientation with respect to RGS000. These three different orientations were chosen to permit ‘decontamination’ of spectra from overlaps with other sources, creating a reconstructed spectrum from successively acquired exposures of a given field with the three red grisms.
Each grism is glued onto six Invar pads, whose coefficient of thermal expansion closely matches silica. The blades are bonded to flexible aluminium-alloy blades that limit stress transmission during the cooldown; these blades are part of the aluminium-alloy mount, which is directly integrated into the grism wheel (Riva et al., 2011).
@@ 3 NISP ground testing setup
After completion of the instrument integration in 2019, the NISP flight model underwent a series of tests in vacuum and with the instrument cooled down to its operational temperature. Two test campaigns were conducted inside the ERIOS vacuum chamber at the ‘Laboratoire d’Astrophysique de Marseille’ in the fall of 2019 and the beginning of 2020, to validate the instrument’s performance and its functionality in a space environment.
The ERIOS chamber (Costille et al., 2016b) is a large cryogenic chamber, with an external envelope of 6 m in length and a diameter of 4 m, capable of achieving high vacuum ( 10-6 mbar) and to cool down an inner volume of 45 m3 to low temperatures ( 80 K) with a high stability (4 mK) thanks to liquid nitrogen shrouds covering the entire inner surface of the chamber.
@@
The entire aperture of the mirror was illuminated through this 2 pin-hole located at the primary focus of the elliptical mirror, creating a point source for NISP. Both, the mirror and the pin-hole were attached to the same baseplate and the whole system (mirror, pin-hole, and baseplate) was made from silica to ensure stability of the mirror’s focal distance. The telescope baseplate was attached to translation and rotation stages using thermal flexures made from a low conductivity material, thermally isolating the telescope, while providing it with 5 degrees of freedom when being pointed towards different positions in NISP’s field-of-view.
@@
Inside ERIOS, light was fed to the telescope through a cryogenic fibre connected to the pin-hole, with light sources being located outside the chamber. At the interface of the ERIOS chamber a pair of vacuum windows connected the fibre with the light sources, using a cooled neutral density filter to suppress thermal background. On the outside, a set of neutral density filters allowed us to select the intensity of light entering the fibre.
@@
Spectral lamps were fed into the monochromator to characterise and calibrate it, using HgCd, He, or Cs lamps for the visible band below , and an Ar lamp for the near-infrared domain. The output of the monochromator was scanned using either silicon or germanium photodiodes. The calibration yielded a selectable maximum wavelength accuracy of , and the monochromator bandwidth was measured to be approximately in Full-Width-Half-Maximum.
The etalon was re-aligned and re-calibrated in warm conditions with a Perkin Elmer Lambda 900 UV/VIS/NIR spectrometer (‘Lambda 900’) before testing of NISP started. During the tests, thermal sensor monitored the stability of the light sources and the etalon. Figure 3 shows the Fabry–Pérot etalon’s spectrum as it was measured and calibrated with a precision of 0.2 nm using the Lambda 900 spectrometer. A comparison to the grisms’s transmission shows the number of Fabry–Pérot’s peaks available for calibration. The grism channels’ total transmission shown here accounts for the transmission of the corrector lens and camera lens assembly lenses, the grisms themselves, as well as the detector quantum efficiency. In this figure neither the Euclid mirrors nor dichroic are accounted for, as both elements were not present in the instrument-level test setup. Note that in Fig. 3, the transmission of the -order components of NISP’s spectrograms, which are below few per cent at maximum, have been multiplied by a factor of 10 to make them stand out. Additionally, the optical transmission shown in Fig. 3 is derived from combined subsystem-level measurements. Due to limited knowledge of the optical ground equipment transmission – including fibre, pinhole, and telescope contribution to the total transmission – and the absence of a calibrated photodiode in the ERIOS chamber, we were unable to measure the absolute transmission of the NISP optics during the ground test campaign, and thus we relied on subsystem measurements to estimate the total transmission presented in Fig. 3.
Prior to the start of the optical tests, the focal length of the NISP instrument was measured at cold (NISP focal plane at and NISP optics at ), using a monochromatic light beam with a wavelength of 1000 nm in the photometric passband. In this process, point spread function widths were measured from spots created by the point source. This was carried out at five positions in the NISP field-of-view, where hundreds of individual slightly dithered exposures were taken to over-sample the PSF and to increase statistical accuracy on the point spread function width measurement. This was repeated for different distances of the telescope simulator from NISP. During the point spread functions acquisitions, the relative distance between NISP and the telescope simulator was accurately monitored with a laser tracker system that measured the position of ten targets attached to the NISP mechanical structure relative to five targets attached to the back of the telescope mirror (Costille et al., 2016a, 2018).
@@
It is important to note that this measurement also established the fixed reference plane for positioning NISP relative to the VIS instrument, ensuring optimal NISP image quality once the Euclid telescope is focused on the VIS instrument, which has more stringent image quality requirements. The relative alignment between NISP and VIS was verified during the Payload Module ground tests and is beyond the scope of this paper. Nevertheless, the excellent optical performance delivered by NISP in flight confirms the accuracy and stability of its alignment with the Euclid telescope.
@@ 4.1 Optical quality of RGS000, RGS180 & BGS000
The NISP optical quality was assessed by evaluating the radii that encircle 50% and 80% of the point spread function’s total energy – in the following these radii will be referred to as EE50 and EE80, respectively. In the spectroscopic channel, this verification was done for every grism of the filter wheel at three monochromatic wavelengths: 1300 nm, 1500 nm, and 1800 nm for all three red grisms and 900 nm, 1120 nm, and 1340 nm for the blue grism. Additionally, we tested image quality of both RGS000 and RGS180 with a grism wheel rotation offset to validate that offsetting the grism positions by this amount would not degrade image quality. This additional measurement was part of an evaluation of a modified survey strategy to overcome the reduced quality of data from RGS270 (see Sect. 4.2).
@@
The image quality was evaluated with an analysis of monochromatic point spread function images taken at the four corners of the NISP field-of-view. Because the NISP plate scale of per pixel undersamples the point spread function, hundreds of point spread function images were acquired around each position, dithering by a tenth of a pixel around the target position in the NISP image plane. This dithering, obtained by moving the telescope on its axes, allowed to spatially over-sample the point spread functions and reduced the statistical uncertainties on the EE50 and EE80 measurements.
@@
Before measuring point spread function properties, images acquired by the NISP instrument were corrected for pixel quantum efficiency and conversion gain, and a NISP residual thermal background acquired under ‘dark test conditions’, i.e. with the NISP filter wheel in closed position and all light sources turned off. An automatic procedure, derived from the work of Jiang et al. (2016), was used to locate the point spread function within the 2k 2k pixels of the target detectors. Once located, a 20 20 pixel stamp centred on the expected point spread function position was extracted from the images and point spread function metrics were extracted by image analyses. Every extracted sub-image was controlled by eye and e.g. detector cosmetic defects misidentified as point spread functions by our algorithm were rejected from the analyses. An Intra Pixel Capacitance correction, relying on a Richardson–Lucy deconvolution method (Lucy, 1974; Richardson, 1972) using Intra Pixel Capacitance coefficients previously measured during detector characterisation as the deconvolution kernel, was applied to each extracted stamp. However, it should be noted that the spread of the measured EE50 and EE80 distributions were much wider than the correction factor applied through this Intra Pixel Capacitance deconvolution. The Intra Pixel Capacitance signal leaking to neighbouring pixels was measured to be very small, of 2.870.01% losses in the central pixel that are non-equally redistributed to the eight immediately adjacent pixels (Le Graët et al., 2022).
@@
Figure 4 shows one example of a single exposure of a NISP point spread function acquisition. The left panel is a full frame image (i.e. 2k 2k pixels) from one detector, and the artificial point source at the top. A ring-like structure is visible surrounding the point spread function, that resulted from scattered light at the exit of the pin-hole. Since the ring radius is very large compared to the point spread function diameter, it was easily disregarded in the analyses. However, these rings were quite useful in validating that the blind search algorithm indeed successfully located the point spread function. In the centre panel of Fig. 4 a zoom into the small box around the point spread function is shown. The diffuse circular feature to the right of the compact point spread function is produced by light emission coming from the fibre core passing through a non-perfectly blocking neutral density filter around the pin-hole. This was meant to block light throughout the full NISP wavelength range. It however appeared that instead it had a transmission increasing with wavelength, starting with an optical density 5 at 900 nm and reaching an optical density of 4.5 at 1800 nm, creating this patch. The rightmost panel shows another zoom step into the point spread function, with contrast adjusted to prevent display saturation.
@@
The EE50 and EE80 are deduced from point spread function model fitting with the following model
@@
where are the spatial coordinates in the pixel mosaic reference frame, is a point spread function model of parameters – for the functions that were considered see below – is a -Bell function (Boyd, 2006) of parameters – which is used to model the contribution from the fibre core – and , a constant introduced to account for a residual constant background. As the fit is limited to a small stamp of 20 20 pixels centred on the point spread function, a constant background is a good first approximation at these scales. Note that the axis is defined such that the light dispersed by RGS000 along a spectral trace has its wavelength increase with .
As shown in Fig. 4, the NISP optical design concentrates most of the photon energy within a single pixel, yielding a point spread function with a Full-Width-Half-Maximum on the order of . Such a small point spread function makes the estimation of EE50 or EE80 particularly challenging. Thus to estimate the reliability of our estimates, we tested three different functions to model the point spread function profile.
@@
An asymmetric-Gaussian function for which amplitude, centroid position, width along the and axes as well as inclination with respect to the axis are free parameters. When using this model, EE50 is evaluated as , with the width of the point spread function in the cross-dispersion direction. We used the cross-dispersion direction to minimise the impact of the monochromator bandwidth on the estimate of EE50. However, such an estimation might be underestimating the true width, by assuming the point spread function is having a Gaussian profile and neglecting potential contributions from a non-Gaussian tail.
@@
The sum of two asymmetric-Gaussian functions, hereafter refereed as a dual-Gaussian profile, for which the respective amplitude, common centroid position, individual width along the and axes, as well as common inclination are free parameters. To avoid any degeneracy in the model fitting, the amplitude and width of the larger Gaussian was defined by a multiplicative factor relative to that of the narrower Gaussian. This model tends to better account for potential non-Gaussian tails in the point spread function profile.